# One Subgoal at a Time: Zero-Shot Generalization to Arbitrary Linear Temporal Logic Requirements in Multi-Task Reinforcement Learning

### v1

@@ 5.1 Can GenZ-LTL zero-shot generalize to arbitrary LTL specifications (Q1)?
LetterWorld $\varphi_{1}$ $0.51_{\pm 0.10}$ $0.00_{\pm 0.00}$ $29.03_{\pm 3.08}$ $0.87_{\pm 0.04}$ $0.02_{\pm 0.02}$ $15.37_{\pm 0.39}$ $0.87_{\pm 0.02}$ $0.00_{\pm 0.00}$ $8.12_{\pm 0.21}$ $0.96_{\pm 0.04}$ $0.00_{\pm 0.00}$ $18.71_{\pm 2.06}$ $0.98_{\pm 0.02}$ $0.00_{\pm 0.00}$ $7.38_{\pm 0.13}$
$\varphi_{2}$ $0.66_{\pm 0.16}$ $0.14_{\pm 0.11}$ $23.25_{\pm 4.93}$ $0.90_{\pm 0.03}$ $0.00_{\pm 0.00}$ $9.26_{\pm 0.75}$ $0.91_{\pm 0.02}$ $0.01_{\pm 0.01}$ $5.88_{\pm 0.12}$ $0.90_{\pm 0.07}$ $0.07_{\pm 0.04}$ $13.64_{\pm 2.71}$ $1.00_{\pm 0.00}$ $0.00_{\pm 0.00}$ $5.48_{\pm 0.09}$
$\varphi_{3}$ $0.75_{\pm 0.07}$ $0.14_{\pm 0.04}$ $24.60_{\pm 4.38}$ $0.72_{\pm 0.06}$ $0.09_{\pm 0.07}$ $14.02_{\pm 1.09}$ $0.81_{\pm 0.02}$ $0.01_{\pm 0.00}$ $9.00_{\pm 0.14}$ $0.86_{\pm 0.05}$ $0.08_{\pm 0.03}$ $20.02_{\pm 2.05}$ $0.96_{\pm 0.01}$ $0.00_{\pm 0.00}$ $8.61_{\pm 0.46}$
$\varphi_{4}$ $0.56_{\pm 0.19}$ $0.00_{\pm 0.00}$ $29.69_{\pm 8.50}$ $0.99_{\pm 0.00}$ $0.00_{\pm 0.00}$ $25.54_{\pm 0.84}$ $0.89_{\pm 0.02}$ $0.00_{\pm 0.00}$ $7.05_{\pm 0.28}$ $0.89_{\pm 0.04}$ $0.00_{\pm 0.00}$ $18.78_{\pm 2.82}$ $0.98_{\pm 0.01}$ $0.00_{\pm 0.00}$ $7.41_{\pm 0.37}$
Ave. $0.62_{\pm 0.16}$ $0.07_{\pm 0.09}$ $26.64_{\pm 5.87}$ $0.87_{\pm 0.11}$ $0.03_{\pm 0.05}$ $16.05_{\pm 6.13}$ $0.87_{\pm 0.04}$ $0.00_{\pm 0.01}$ $7.51_{\pm 1.21}$ $0.90_{\pm 0.06}$ $0.04_{\pm 0.04}$ $17.79_{\pm 3.37}$ $0.98_{\pm 0.02}$ $0.00_{\pm 0.00}$ $7.22_{\pm 1.18}$
$\varphi_{5}$ $0.30_{\pm 0.13}$ $0.00_{\pm 0.00}$ $39.12_{\pm 2.08}$ $0.67_{\pm 0.07}$ $0.07_{\pm 0.05}$ $20.53_{\pm 0.49}$ $0.78_{\pm 0.05}$ $0.00_{\pm 0.00}$ $11.24_{\pm 0.39}$ $0.88_{\pm 0.05}$ $0.00_{\pm 0.00}$ $26.02_{\pm 1.62}$ $0.99_{\pm 0.00}$ $0.00_{\pm 0.00}$ $10.60_{\pm 0.12}$
$\varphi_{6}$ $0.18_{\pm 0.10}$ $0.00_{\pm 0.00}$ $28.40_{\pm 6.15}$ $0.70_{\pm 0.08}$ $0.15_{\pm 0.05}$ $14.69_{\pm 0.75}$ $0.81_{\pm 0.02}$ $0.00_{\pm 0.00}$ $8.23_{\pm 0.18}$ $0.90_{\pm 0.04}$ $0.00_{\pm 0.00}$ $19.48_{\pm 1.96}$ $0.95_{\pm 0.01}$ $0.00_{\pm 0.00}$ $7.48_{\pm 0.06}$
$\varphi_{7}$ $0.29_{\pm 0.16}$ $0.33_{\pm 0.20}$ $40.71_{\pm 1.43}$ $0.59_{\pm 0.05}$ $0.14_{\pm 0.07}$ $19.52_{\pm 1.25}$ $0.71_{\pm 0.02}$ $0.03_{\pm 0.01}$ $11.82_{\pm 0.39}$ $0.73_{\pm 0.05}$ $0.17_{\pm 0.05}$ $26.07_{\pm 3.73}$ $0.94_{\pm 0.01}$ $0.00_{\pm 0.00}$ $10.97_{\pm 0.08}$
$\varphi_{8}$ $0.20_{\pm 0.08}$ $0.47_{\pm 0.22}$ $37.50_{\pm 4.16}$ $0.63_{\pm 0.09}$ $0.09_{\pm 0.07}$ $20.10_{\pm 0.82}$ $0.75_{\pm 0.02}$ $0.02_{\pm 0.01}$ $11.20_{\pm 0.25}$ $0.76_{\pm 0.11}$ $0.13_{\pm 0.05}$ $25.60_{\pm 2.89}$ $0.93_{\pm 0.01}$ $0.00_{\pm 0.00}$ $10.24_{\pm 0.08}$
Ave. $0.24_{\pm 0.12}$ $0.20_{\pm 0.25}$ $36.43_{\pm 6.08}$ $0.65_{\pm 0.08}$ $0.11_{\pm 0.06}$ $18.71_{\pm 2.54}$ $0.76_{\pm 0.05}$ $0.01_{\pm 0.02}$ $10.62_{\pm 1.47}$ $0.82_{\pm 0.10}$ $0.07_{\pm 0.08}$ $24.29_{\pm 3.77}$ $0.95_{\pm 0.03}$ $0.00_{\pm 0.00}$ $9.82_{\pm 1.42}$
ZoneEnv $\varphi_{9}$ $0.13_{\pm 0.08}$ $0.18_{\pm 0.25}$ $496.09_{\pm 128.52}$ $0.88_{\pm 0.05}$ $0.03_{\pm 0.00}$ $474.41_{\pm 23.42}$ $0.89_{\pm 0.03}$ $0.04_{\pm 0.00}$ $327.80_{\pm 23.78}$ $0.90_{\pm 0.07}$ $0.08_{\pm 0.07}$ $423.11_{\pm 51.08}$ $0.99_{\pm 0.01}$ $0.01_{\pm 0.01}$ $380.15_{\pm 5.17}$
$\varphi_{10}$ $0.39_{\pm 0.32}$ $0.38_{\pm 0.24}$ $421.46_{\pm 58.76}$ $0.88_{\pm 0.03}$ $0.06_{\pm 0.01}$ $303.43_{\pm 4.46}$ $0.88_{\pm 0.03}$ $0.09_{\pm 0.01}$ $221.94_{\pm 23.26}$ $0.92_{\pm 0.02}$ $0.06_{\pm 0.02}$ $301.21_{\pm 101.99}$ $0.97_{\pm 0.01}$ $0.02_{\pm 0.02}$ $253.44_{\pm 10.22}$
$\varphi_{11}$ $0.82_{\pm 0.15}$ $0.08_{\pm 0.03}$ $284.41_{\pm 95.62}$ $0.86_{\pm 0.04}$ $0.06_{\pm 0.03}$ $306.94_{\pm 11.26}$ $0.91_{\pm 0.06}$ $0.04_{\pm 0.02}$ $215.42_{\pm 15.01}$ $0.93_{\pm 0.06}$ $0.04_{\pm 0.02}$ $247.83_{\pm 42.98}$ $0.99_{\pm 0.01}$ $0.01_{\pm 0.01}$ $249.57_{\pm 11.85}$
$\varphi_{12}$ $0.94_{\pm 0.03}$ $0.02_{\pm 0.03}$ $122.89_{\pm 20.97}$ $0.90_{\pm 0.01}$ $0.05_{\pm 0.02}$ $136.32_{\pm 10.27}$ $0.97_{\pm 0.01}$ $0.01_{\pm 0.00}$ $116.40_{\pm 13.07}$ $1.00_{\pm 0.01}$ $0.00_{\pm 0.00}$ $105.97_{\pm 17.31}$ $1.00_{\pm 0.00}$ $0.00_{\pm 0.00}$ $135.60_{\pm 8.06}$
Ave. $0.57_{\pm 0.37}$ $0.17_{\pm 0.21}$ $331.21_{\pm 165.88}$ $0.88_{\pm 0.04}$ $0.05_{\pm 0.02}$ $305.28_{\pm 123.33}$ $0.91_{\pm 0.05}$ $0.04_{\pm 0.03}$ $220.39_{\pm 78.77}$ $0.94_{\pm 0.05}$ $0.04_{\pm 0.05}$ $269.53_{\pm 129.95}$ $0.99_{\pm 0.01}$ $0.01_{\pm 0.01}$ $254.69_{\pm 89.18}$
$\varphi_{13}$ $0.17_{\pm 0.23}$ $0.01_{\pm 0.01}$ $886.66_{\pm 104.03}$ $0.73_{\pm 0.06}$ $0.09_{\pm 0.02}$ $618.42_{\pm 22.80}$ $0.91_{\pm 0.03}$ $0.00_{\pm 0.00}$ $508.26_{\pm 54.81}$ $0.69_{\pm 0.17}$ $0.00_{\pm 0.00}$ $664.59_{\pm 29.91}$ $0.98_{\pm 0.02}$ $0.00_{\pm 0.00}$ $436.25_{\pm 18.35}$
$\varphi_{14}$ $0.36_{\pm 0.29}$ $0.01_{\pm 0.01}$ $885.13_{\pm 148.93}$ $0.68_{\pm 0.06}$ $0.11_{\pm 0.06}$ $611.24_{\pm 38.37}$ $0.91_{\pm 0.03}$ $0.00_{\pm 0.00}$ $473.45_{\pm 52.67}$ $0.77_{\pm 0.21}$ $0.00_{\pm 0.00}$ $628.57_{\pm 22.89}$ $0.98_{\pm 0.01}$ $0.00_{\pm 0.00}$ $434.45_{\pm 14.61}$
$\varphi_{15}$ $0.08_{\pm 0.05}$ $0.49_{\pm 0.30}$ $855.56_{\pm 311.81}$ $0.70_{\pm 0.06}$ $0.10_{\pm 0.03}$ $599.71_{\pm 18.62}$ $0.78_{\pm 0.05}$ $0.12_{\pm 0.05}$ $523.50_{\pm 64.73}$ $0.62_{\pm 0.09}$ $0.02_{\pm 0.01}$ $680.11_{\pm 9.51}$ $0.97_{\pm 0.01}$ $0.01_{\pm 0.00}$ $388.01_{\pm 15.47}$
$\varphi_{16}$ $0.03_{\pm 0.04}$ $0.02_{\pm 0.04}$ $931.19_{\pm 420.81}$ $0.68_{\pm 0.09}$ $0.08_{\pm 0.04}$ $596.29_{\pm 25.87}$ $0.89_{\pm 0.10}$ $0.00_{\pm 0.00}$ $515.48_{\pm 45.53}$ $0.70_{\pm 0.18}$ $0.00_{\pm 0.00}$ $616.28_{\pm 59.88}$ $0.99_{\pm 0.01}$ $0.00_{\pm 0.00}$ $396.83_{\pm 20.39}$
Ave. $0.12_{\pm 0.18}$ $0.17_{\pm 0.28}$ $886.39_{\pm 288.60}$ $0.70_{\pm 0.07}$ $0.09_{\pm 0.04}$ $606.42_{\pm 26.76}$ $0.87_{\pm 0.08}$ $0.03_{\pm 0.06}$ $505.17_{\pm 54.03}$ $0.70_{\pm 0.15}$ $0.00_{\pm 0.01}$ $647.39_{\pm 40.74}$ $0.98_{\pm 0.02}$ $0.01_{\pm 0.01}$ $408.71_{\pm 27.47}$
@@
Finite-horizon tasks. We first evaluate finite-horizon tasks, i.e. tasks that can be satisfied with finite-length trajectories. All the specifications considered are complex LTL formulas with nested temporal and Boolean operators (details in Table 3). The evaluation results are shown in Table 1. Our method consistently outperforms the baselines in terms of both success and violation rates. LTL2Action encodes the full structure of the current LTL specification, making it less effective at adapting to out-of-distribution (OOD) specifications at test time. GCRL-LTL learns a goal-conditioned policy but does not model safety constraints explicitly, limiting its ability to enforce constraint satisfaction. DeepLTL relies on curriculum training which samples random reach-avoid subgoal sequences only up to a certain length. When evaluating specifications $\varphi_{9}$$\varphi_{12}$, although the specifications themselves are unseen, the underlying subgoal sequences are likely to have been encountered during training, leading to better efficiency, i.e., smaller $\mu$. However, for specifications with longer subgoal sequences, e.g, $\varphi_{13}$$\varphi_{16}$, which are OOD with respect to the training sequences, DeepLTL exhibits reduced performance. RAD-embeddings pre-trains the automaton embeddings by sampling from a fixed distribution. As a result, it also struggles with the issue of OOD generalization when faced with unseen automata, which is reflected in its performance degradation under more complex LTL specifications: the success rate $\eta_{s}$ drops from 0.9 on $\varphi_{1-4}$ to 0.82 on $\varphi_{5-8}$ and 0.94 on $\varphi_{9-12}$ to 0.7 on $\varphi_{13-16}$. In contrast, our method conditions only on the current subgoal, sidestepping the issue of OOD subgoal sequences. By incorporating an HJ reachability constraint, our method improves adherence to safety constraints. Additionally, since we optimize for the discounted cumulative reward obtained upon satisfying a subgoal, efficient policies that take fewer steps are encouraged.
@@
GenZ-LTL(ours) $\mu_{\text{acc}}\uparrow$ $208.95_{\pm 14.39}$ $102.12_{\pm 7.02}$ $55.17_{\pm 1.08}$ $55.16_{\pm 4.23}$ $32.75_{\pm 1.20}$ $8135.67_{\pm 1489.99}$
$\eta_{v}\downarrow$ $0.00_{\pm 0.00}$ $0.00_{\pm 0.00}$ $0.00_{\pm 0.00}$ $0.07_{\pm 0.02}$ $0.03_{\pm 0.01}$ $0.03_{\pm 0.02}$

### v6

@@ 5.1 Can GenZ-LTL zero-shot generalize to arbitrary LTL specifications (Q1)?
LetterWorld $\varphi_{1}$ $51_{\pm 0.10} 0.0$$00_{\pm 0.00}$ $29.03_{\pm 3.08}$ $87_{\pm 0.04} 0.0$$02_{\pm 0.02}$ $15.37_{\pm 0.39}$ $87_{\pm 0.02} 0.0$$00_{\pm 0.00}$ $8.12_{\pm 0.21}$ $96_{\pm 0.04} 0.0$$00_{\pm 0.00}$ $18.71_{\pm 2.06}$ $98_{\pm 0.02} 0.0$$00_{\pm 0.00}$ $7.38_{\pm 0.13}$
$\varphi_{2}$ $0.66_{\pm 0.16}$ $0.14_{\pm 0.11}$ $23.25_{\pm 4.93}$ $90_{\pm 0.03} 0.0$$00_{\pm 0.00}$ $9.26_{\pm 0.75}$ $91_{\pm 0.02} 0.0$$01_{\pm 0.01}$ $5.88_{\pm 0.12}$ $90_{\pm 0.07} 0.0$$07_{\pm 0.04}$ $13.64_{\pm 2.71}$ $1.00_{\pm 0.00}$ $0.00_{\pm 0.00}$ $5.48_{\pm 0.09}$
$\varphi_{3}$ $0.75_{\pm 0.07}$ $0.14_{\pm 0.04}$ $24.60_{\pm 4.38}$ $72_{\pm 0.06} 0.0$$09_{\pm 0.07}$ $14.02_{\pm 1.09}$ $81_{\pm 0.02} 0.0$$01_{\pm 0.00}$ $9.00_{\pm 0.14}$ $86_{\pm 0.05} 0.0$$08_{\pm 0.03}$ $20.02_{\pm 2.05}$ $96_{\pm 0.01} 0.0$$00_{\pm 0.00}$ $8.61_{\pm 0.46}$
$\varphi_{4}$ $56_{\pm 0.19} 0.0$$00_{\pm 0.00}$ $29.69_{\pm 8.50}$ $99_{\pm 0.00} 0.0$$00_{\pm 0.00}$ $25.54_{\pm 0.84}$ $89_{\pm 0.02} 0.0$$00_{\pm 0.00}$ $7.05_{\pm 0.28}$ $89_{\pm 0.04} 0.0$$00_{\pm 0.00}$ $18.78_{\pm 2.82}$ $98_{\pm 0.01} 0.0$$00_{\pm 0.00}$ $7.41_{\pm 0.37}$
Ave. $62_{\pm 0.16} 0.0$$07_{\pm 0.09}$ $26.64_{\pm 5.87}$ $87_{\pm 0.11} 0.0$$03_{\pm 0.05}$ $16.05_{\pm 6.13}$ $87_{\pm 0.04} 0.0$$00_{\pm 0.01}$ $7.51_{\pm 1.21}$ $90_{\pm 0.06} 0.0$$04_{\pm 0.04}$ $17.79_{\pm 3.37}$ $98_{\pm 0.02} 0.0$$00_{\pm 0.00}$ $7.22_{\pm 1.18}$
$\varphi_{5}$ $30_{\pm 0.13} 0.0$$00_{\pm 0.00}$ $39.12_{\pm 2.08}$ $67_{\pm 0.07} 0.0$$07_{\pm 0.05}$ $20.53_{\pm 0.49}$ $78_{\pm 0.05} 0.0$$00_{\pm 0.00}$ $11.24_{\pm 0.39}$ $88_{\pm 0.05} 0.0$$00_{\pm 0.00}$ $26.02_{\pm 1.62}$ $99_{\pm 0.00} 0.0$$00_{\pm 0.00}$ $10.60_{\pm 0.12}$
$\varphi_{6}$ $18_{\pm 0.10} 0.0$$00_{\pm 0.00}$ $28.40_{\pm 6.15}$ $0.70_{\pm 0.08}$ $0.15_{\pm 0.05}$ $14.69_{\pm 0.75}$ $81_{\pm 0.02} 0.0$$00_{\pm 0.00}$ $8.23_{\pm 0.18}$ $90_{\pm 0.04} 0.0$$00_{\pm 0.00}$ $19.48_{\pm 1.96}$ $95_{\pm 0.01} 0.0$$00_{\pm 0.00}$ $7.48_{\pm 0.06}$
$\varphi_{7}$ $0.29_{\pm 0.16}$ $0.33_{\pm 0.20}$ $40.71_{\pm 1.43}$ $0.59_{\pm 0.05}$ $0.14_{\pm 0.07}$ $19.52_{\pm 1.25}$ $71_{\pm 0.02} 0.0$$03_{\pm 0.01}$ $11.82_{\pm 0.39}$ $0.73_{\pm 0.05}$ $0.17_{\pm 0.05}$ $26.07_{\pm 3.73}$ $94_{\pm 0.01} 0.0$$00_{\pm 0.00}$ $10.97_{\pm 0.08}$
$\varphi_{8}$ $0.20_{\pm 0.08}$ $0.47_{\pm 0.22}$ $37.50_{\pm 4.16}$ $63_{\pm 0.09} 0.0$$09_{\pm 0.07}$ $20.10_{\pm 0.82}$ $75_{\pm 0.02} 0.0$$02_{\pm 0.01}$ $11.20_{\pm 0.25}$ $0.76_{\pm 0.11}$ $0.13_{\pm 0.05}$ $25.60_{\pm 2.89}$ $93_{\pm 0.01} 0.0$$00_{\pm 0.00}$ $10.24_{\pm 0.08}$
Ave. $0.24_{\pm 0.12}$ $0.20_{\pm 0.25}$ $36.43_{\pm 6.08}$ $0.65_{\pm 0.08}$ $0.11_{\pm 0.06}$ $18.71_{\pm 2.54}$ $76_{\pm 0.05} 0.0$$01_{\pm 0.02}$ $10.62_{\pm 1.47}$ $82_{\pm 0.10} 0.0$$07_{\pm 0.08}$ $24.29_{\pm 3.77}$ $95_{\pm 0.03} 0.0$$00_{\pm 0.00}$ $9.82_{\pm 1.42}$
ZoneEnv $\varphi_{9}$ $0.13_{\pm 0.08}$ $0.18_{\pm 0.25}$ $496.09_{\pm 128.52}$ $88_{\pm 0.05} 0.0$$03_{\pm 0.00}$ $474.41_{\pm 23.42}$ $89_{\pm 0.03} 0.0$$04_{\pm 0.00}$ $327.80_{\pm 23.78}$ $90_{\pm 0.07} 0.0$$08_{\pm 0.07}$ $423.11_{\pm 51.08}$ $99_{\pm 0.01} 0.0$$01_{\pm 0.01}$ $380.15_{\pm 5.17}$
$\varphi_{10}$ $0.39_{\pm 0.32}$ $0.38_{\pm 0.24}$ $421.46_{\pm 58.76}$ $88_{\pm 0.03} 0.0$$06_{\pm 0.01}$ $303.43_{\pm 4.46}$ $88_{\pm 0.03} 0.0$$09_{\pm 0.01}$ $221.94_{\pm 23.26}$ $92_{\pm 0.02} 0.0$$06_{\pm 0.02}$ $301.21_{\pm 101.99}$ $97_{\pm 0.01} 0.0$$02_{\pm 0.02}$ $253.44_{\pm 10.22}$
$\varphi_{11}$ $82_{\pm 0.15} 0.0$$08_{\pm 0.03}$ $284.41_{\pm 95.62}$ $86_{\pm 0.04} 0.0$$06_{\pm 0.03}$ $306.94_{\pm 11.26}$ $91_{\pm 0.06} 0.0$$04_{\pm 0.02}$ $215.42_{\pm 15.01}$ $93_{\pm 0.06} 0.0$$04_{\pm 0.02}$ $247.83_{\pm 42.98}$ $99_{\pm 0.01} 0.0$$01_{\pm 0.01}$ $249.57_{\pm 11.85}$
$\varphi_{12}$ $94_{\pm 0.03} 0.0$$02_{\pm 0.03}$ $122.89_{\pm 20.97}$ $90_{\pm 0.01} 0.0$$05_{\pm 0.02}$ $136.32_{\pm 10.27}$ $97_{\pm 0.01} 0.0$$01_{\pm 0.00}$ $116.40_{\pm 13.07}$ $1.00_{\pm 0.01}$ $0.00_{\pm 0.00}$ $105.97_{\pm 17.31}$ $1.00_{\pm 0.00}$ $0.00_{\pm 0.00}$ $135.60_{\pm 8.06}$
Ave. $0.57_{\pm 0.37}$ $0.17_{\pm 0.21}$ $331.21_{\pm 165.88}$ $88_{\pm 0.04} 0.0$$05_{\pm 0.02}$ $305.28_{\pm 123.33}$ $91_{\pm 0.05} 0.0$$04_{\pm 0.03}$ $220.39_{\pm 78.77}$ $94_{\pm 0.05} 0.0$$04_{\pm 0.05}$ $269.53_{\pm 129.95}$ $99_{\pm 0.01} 0.0$$01_{\pm 0.01}$ $254.69_{\pm 89.18}$
$\varphi_{13}$ $17_{\pm 0.23} 0.0$$01_{\pm 0.01}$ $886.66_{\pm 104.03}$ $73_{\pm 0.06} 0.0$$09_{\pm 0.02}$ $618.42_{\pm 22.80}$ $91_{\pm 0.03} 0.0$$00_{\pm 0.00}$ $508.26_{\pm 54.81}$ $69_{\pm 0.17} 0.0$$00_{\pm 0.00}$ $664.59_{\pm 29.91}$ $98_{\pm 0.02} 0.0$$00_{\pm 0.00}$ $436.25_{\pm 18.35}$
$\varphi_{14}$ $36_{\pm 0.29} 0.0$$01_{\pm 0.01}$ $885.13_{\pm 148.93}$ $0.68_{\pm 0.06}$ $0.11_{\pm 0.06}$ $611.24_{\pm 38.37}$ $91_{\pm 0.03} 0.0$$00_{\pm 0.00}$ $473.45_{\pm 52.67}$ $77_{\pm 0.21} 0.0$$00_{\pm 0.00}$ $628.57_{\pm 22.89}$ $98_{\pm 0.01} 0.0$$00_{\pm 0.00}$ $434.45_{\pm 14.61}$
$\varphi_{15}$ $0.08_{\pm 0.05}$ $0.49_{\pm 0.30}$ $855.56_{\pm 311.81}$ $0.70_{\pm 0.06}$ $0.10_{\pm 0.03}$ $599.71_{\pm 18.62}$ $0.78_{\pm 0.05}$ $0.12_{\pm 0.05}$ $523.50_{\pm 64.73}$ $62_{\pm 0.09} 0.0$$02_{\pm 0.01}$ $680.11_{\pm 9.51}$ $97_{\pm 0.01} 0.0$$01_{\pm 0.00}$ $388.01_{\pm 15.47}$
$\varphi_{16}$ $03_{\pm 0.04} 0.0$$02_{\pm 0.04}$ $931.19_{\pm 420.81}$ $68_{\pm 0.09} 0.0$$08_{\pm 0.04}$ $596.29_{\pm 25.87}$ $89_{\pm 0.10} 0.0$$00_{\pm 0.00}$ $515.48_{\pm 45.53}$ $70_{\pm 0.18} 0.0$$00_{\pm 0.00}$ $616.28_{\pm 59.88}$ $99_{\pm 0.01} 0.0$$00_{\pm 0.00}$ $396.83_{\pm 20.39}$
Ave. $0.12_{\pm 0.18}$ $0.17_{\pm 0.28}$ $886.39_{\pm 288.60}$ $70_{\pm 0.07} 0.0$$09_{\pm 0.04}$ $606.42_{\pm 26.76}$ $87_{\pm 0.08} 0.0$$03_{\pm 0.06}$ $505.17_{\pm 54.03}$ $70_{\pm 0.15} 0.0$$00_{\pm 0.01}$ $647.39_{\pm 40.74}$ $98_{\pm 0.02} 0.0$$01_{\pm 0.01}$ $408.71_{\pm 27.47}$
@@
Finite-horizon tasks. We first evaluate finite-horizon tasks, i.e. tasks that can be satisfied with finite-length trajectories. All the specifications considered are complex LTL formulas with nested temporal and Boolean operators (details in Table 1). The evaluation results are shown in Table 2. Our method consistently outperforms the baselines in terms of both success and violation rates. LTL2Action encodes the full structure of the current LTL specification, making it less effective at adapting to out-of-distribution (OOD) specifications at test time. GCRL-LTL learns a goal-conditioned policy but does not model safety constraints explicitly, limiting its ability to enforce constraint satisfaction. DeepLTL relies on curriculum training which samples random reach-avoid subgoal sequences only up to a certain length. When evaluating specifications $\varphi_{9}$$\varphi_{12}$, although the specifications themselves are unseen, the underlying subgoal sequences are likely to have been encountered during training, leading to better efficiency, i.e., smaller $\mu$. However, for specifications with longer subgoal sequences, e.g, $\varphi_{13}$$\varphi_{16}$, which are OOD with respect to the training sequences, DeepLTL exhibits reduced performance. RAD-embeddings pre-trains the automaton embeddings by sampling from a fixed distribution. As a result, it also struggles with the issue of OOD generalization when faced with unseen automata, which is reflected in its performance degradation under more complex LTL specifications: the success rate $\eta_{s}$ drops from 0.9 on $\varphi_{1-4}$ to 0.82 on $\varphi_{5-8}$ and 0.94 on $\varphi_{9-12}$ to 0.7 on $\varphi_{13-16}$. In contrast, our method conditions only on the current subgoal, sidestepping the issue of OOD subgoal sequences. By incorporating an HJ reachability constraint, our method improves adherence to safety constraints. Additionally, since we optimize for the discounted cumulative reward obtained upon satisfying a subgoal, efficient policies that take fewer steps are encouraged. We further evaluate our method in ZoneEnv with agents featuring more complex dynamics, which induce additional hard constraints such as preventing head-down falls. Our method consistently outperforms the baselines in both success and violation rates. Details of the setup and results are provided in Appendix C.
@@
GenZ-LTL(ours) $\mu_{\text{acc}}\uparrow$ $208.95_{\pm 14.39}$ $102.12_{\pm 7.02}$ $5.17_{\pm 1.08} 55$$5.16_{\pm 4.23}$ $32.75_{\pm 1.20}$ $8135.67_{\pm 1489.99}$
$\eta_{v}\downarrow$ $_{\pm 0.00} 0.00±0.00$$_{\pm 0.00}$ $00_{\pm 0.00} 0.0$$7_{\pm 0.02} 0.0$$3_{\pm 0.01} 0.0$$03_{\pm 0.02}$